# Near-Optimal Decision Trees in a SPLIT Second

### v1

@@ 1 Introduction
However, these methods tend to yield suboptimal results, lacking general guarantees on either sparsity or accuracy. Recent advances in decision tree algorithms use dynamic programming techniques combined with branch-and-bound strategies, offering solutions that are faster than brute-force approaches and provably optimal (Lin et al., 2020; Aglin et al., 2020; Demirović et al., 2022; McTavish et al., 2022). In fact, Demirović et al. (2022) and (van der Linden et al., 2024) reveal an average gap of $1$-$2$ percentage points between greedy and optimal trees, with Demirović et al. (2022) showing that some datasets can exhibit gaps as large as $10$ percentage points. These algorithms struggle to scale to datasets with hundreds or thousands of features or to deeper trees. It seems that we should return to greedy methods for larger-scale problems, but this would come at a loss of performance. Ideally, we should leverage greed only when it does not significantly deviate from optimality and use dynamic programming otherwise. Dynamic programming approaches build trees recursively, downward from the root. Problems farther from the root contain fewer samples and produce fewer splits. As we show, greedy splits near the root sacrifice performance, while greedy splits near the leaves produce performance close to the optimal. This suggests that we can tolerate less precision on problems close to leaves than on problems closer to the root – and that full optimization on those problems closer to the leaves yields only marginal returns relative to greedy, since we only have a few splits remaining. This has enormous implications, since the number of candidate trees increases exponentially with increases in depth; using greedy splitting closer to the leaves of the tree massively reduces the search space.
@@
Our contributions are as follows. - We develop a family of decision tree algorithms that scale with the dataset size and number of features comparably to standard greedy algorithms but produce trees that are as accurate and sparse as optimal ones (e.g., Lin et al., 2020).
@@ 2 Related Work
Among the many methods for globally optimizing trees, Branch-and-bound approaches with dynamic programming are state of the art for scalability, because they exploit the structure of decision trees (Costa & Pedreira, 2023; Lin et al., 2020; Demirović et al., 2022; McTavish et al., 2022; Aglin et al., 2020). While many other methods exist for optimizing trees, such as MIP solvers (Bertsimas & Dunn, 2017; Verwer & Zhang, 2019), we focus our discussion and comparison of globally optimal decision tree methods on the currently fastest types of approaches - branch and bound with dynamic programming. These approaches search through the space of decision trees while tracking lower and upper bounds of the overall objective at each split to reduce the search space. They can find optimal trees on medium-sized datasets with tens of features and shallow maximum tree depths (Sullivan et al., 2024; Aglin et al., 2020; Lin et al., 2020; Demirović et al., 2022). Aglin et al. (2020) uses a dynamic programming with branch and bound (DPBnB) method that uses advanced caching techniques to find optimal decision trees, though it does not explicitly optimize for sparsity. In contrast, Lin et al. (2020); Hu et al. (2019) use a DPBnB approach to find a tree that optimizes a weighted combination of empirical risk and sparsity, defined by the number of leaves in the tree. McTavish et al. (2022) further enhances this approach by incorporating smart guessing strategies to construct tighter lower bounds for DPBnB, resulting in computational speedups. Demirović et al. (2022) takes a different approach by focusing on finding the optimal tree with a hard constraint on the number of permissible nodes, using advanced caching techniques and an optimized depth-2 decision tree solver. Mazumder et al. (2022) addresses continuous features by defining lower and upper bounds based on quantiles of feature distributions. However, their method is applicable only to shallow optimal trees with depth $\leq 3$, limiting its utility in scenarios with higher-order feature interactions.
@@
Several other approaches are compatible with branch and bound techniques. Blanc et al. (2024) seeks to bridge the gap between greedy and optimal decision trees by selecting a fixed subset of the top $k$ feature splits for each sub-problem. However, this framework does not explicitly account for sparsity. Further, the method is limited by using a global setting for search precision: the approach considers the same number of candidate splits at each subproblem. As we show in our experiments, there is merit to tailoring the level of search precision to parts of the search space where it is most needed.
There are a few methods that use probabilistic search techniques to optimize trees. Sullivan et al. (2024) takes a Bayesian approach, finding the maximum-a-posteriori tree by optimizing over an AND/OR graph, akin to the graph used in earlier branch-and-bound methods like that of Lin et al. (2020). Although their method demonstrates strong performance, their experimental results reveal that it is not responsive to sparsity-inducing hyperparameters – accordingly, we found in our experiments that the method struggles to optimize for sparsity.
@@ 3 Preliminaries
In the Appendix Section A.2, we use Rashomon sets to investigate properties of near-optimal trees.
@@ 4.1 SParse Lookahead for Interpretable Trees (SPLIT)
$\begin{cases}\lambda+\min\Bigg{\{}\frac{|D^{-}|}{|D|},\frac{|D^{+}|}{|D|}\Bigg{\}}\ \ \hskip 44.10185pt\text{if $d^{\prime}=0$}\\ \\ \begin{aligned} &\lambda+\min\Bigg{\{}\frac{|D^{-}|}{|D|},\frac{|D^{+}|}{|D|},\\ &\min_{f\in\mathcal{F}}\Big{\{}L\Big{(}T_{g}\big{(}D(f),d^{\prime},\lambda\big{)}\Big{)}+L\Big{(}T_{g}\big{(}D(\bar{f}),d^{\prime},\lambda\big{)}\Big{)}\Big{\}}\Bigg{\}}\end{aligned}\ \ \\ \hskip 142.26378pt\text{if $d^{\prime}=d-d_{l}$}\\ \begin{aligned} &\lambda+\min\Bigg{\{}\frac{|D^{-}|}{|D|},\frac{|D^{+}|}{|D|},\\ &\min_{f\in\mathcal{F}}\Big{\{}\mathcal{L}\Big{(}D(f),d^{\prime}-1,\lambda\Big{)}+\mathcal{L}\Big{(}D(\bar{f}),d^{\prime}-1,\lambda\Big{)}\Big{\}}\Bigg{\}}\end{aligned}\ \ \\ \hskip 142.26378pt\text{if $d^{\prime}>d-d_{l}$.}\end{cases}$ (3)
@@ 5 Theoretical Analysis of Runtime and Optimality
Corollaries 5.2 and 5.3 show that, compared to globally optimal approaches, we see substantial improvements in runtime when lookahead depth is around half the global search depth.
@@ 6 Experiments
Figure 3: A comparison between the performance of our algorithms and their contemporaries. The lower plots are zoomed-in versions of the upper plots, but with the $x$-axis now being the runtime. SPLIT and LicketySPLIT consistently lie on the bottom left of the test loss-sparsity frontier, with runtimes orders of magnitude faster than many contemporaries. Our algorithms also offer the ideal compromise between runtime and loss. All metrics are averaged over $3$ test-train splits.
@@ A.1.2 What about depth $4$ trees?
In this section, we perform the same evaluation as above, but with depth $4$ trees. We set the lookahead depth as $1$.
Figure 5: A performance comparison between our algorithm and those in literature. Depth $4$ Lookahead depth $1$
@@ A.1.3 What about depth $6$ trees?
Figure 6: A performance comparison between our algorithm and those in literature. Depth $6$ Lookahead depth $2$
A.2 Motivation for SPLIT: Near Optimal Trees Are Generally Greedier Near the Leaves
To test the claim that near optimal trees are likely to behave greedily closer to the leaf, we first generate the Rashomon set of decision trees (Xin et al., 2022) for various values of sparsity penalty $\lambda$ and Rashomon bound $\epsilon$. Let $T\in\mathcal{R}(D,\lambda,\epsilon,d)$ be a tree in the Rashomon set, and let $n\in T$ be any node in $T$. Then, we compute the fraction of all nodes at a given level $\ell\leq d$ (where level 00 corresponds to the root) that were greedy (by which we mean that the split at this node in the tree is optimal with respect to information gain). This corresponds to the following proportion:
$\frac{\sum\limits_{T\in\mathcal{R}(D,\lambda,\epsilon,d)}\sum\limits_{\textrm{$n$}\in T}\mathbbm{1}[\textrm{$n$ is greedy $\land$ level$(n)=\ell$}]}{\sum\limits_{T\in\mathcal{R}(D,\lambda,\epsilon,d)}\sum\limits_{\textrm{$n$}\in T}\mathbbm{1}[\textrm{level}(n)=\ell]}$ (4)
Figure 7 shows the results of this enumeration for $6$ different datasets for different values of $\epsilon$ and $\lambda$. We note that there is a general increase in percentage of greedy splits as one goes deeper in the tree.
Figure 7: A heatmap of the proportion of splits of trees in the Rashomon set that are greedy, stratified by level, for different ($\lambda,\epsilon$) combinations. Only $4$ levels are shown as the $5^{th}$ level corresponds to the leaf. The greyed out regions in the bottom right of a plot represent ($\lambda,\epsilon$) for which the Rashomon set did not contain any trees of that depth. Generally, as you approach the leaves, the proportion of splits appearing in $\epsilon$-optimal trees become increasingly greedy. This is especially noticeable for the Netherlands, Covertype, and COMPAS datasets.
A.3 Many Near-Optimal Trees Exhibit Monotonically Decreasing Optimality Gaps Closer to Leaves
@@
$\beta(T,D,l)=\frac{\sum\limits_{\textrm{$t$}\in T}\delta(D_{t},t)\mathbbm{1}[\textrm{$t$ is rooted at level $\ell$}]}{{\sum\limits_{\textrm{$t$}\in T}\mathbbm{1}[\textrm{$t$ is rooted at level $\ell$}]}}.$ (6)
@@
A.4 Miscellaneous Properties of SPLIT
A.4.1 Which Lookahead Depth Should I Use?
@@
A.4.2 Are SPLIT Trees in the Rashomon Set?
@@
A.4.3 SPLIT with Optimality Preserving Discretization
@@
A.4.4 What is the performance gap between GOSDT post-processing for SPLIT / LicketySPLIT and purely greedy post-processing?
@@
A.5 SPLIT and LicketySPLIT Scaling Experiments
@@
A.6 Rashomon Importance Distribution Under RESPLIT vs TreeFARMS: Threshold Guessing
@@
A.7 Rashomon Importance Distribution Under RESPLIT: Quantile Binarization
@@
A.8 Experimental Setup
A.8.1 Datasets
@@ A.8.2 Preprocessing
    • For version $1$, we chose $n_{est}$ for each dataset such that the resulting binarized dataset has between $40$-$100$ features. This is the version used for experiments in Figures 3, 4, 5, 6 when we compare SPLIT / LicketySPLIT with other datasets.
@@
A.8.3 Details of Comparative Experiments for SPLIT and LicketySPLIT
@@
A.8.4 Description of Machines Used

### v2

@@ 1 Introduction
However, these methods tend to yield suboptimal results, lacking general guarantees on either sparsity or accuracy. Recent advances in decision tree algorithms use dynamic programming techniques combined with branch-and-bound strategies, offering solutions that are faster than brute-force approaches and provably optimal (Lin et al., 2020; Aglin et al., 2020; Demirović et al., 2022; McTavish et al., 2022). In fact, Demirović et al. (2022) and van der Linden et al. (2024) reveal an average gap of $1$-$2$ percentage points between greedy and optimal trees, with Demirović et al. (2022) showing that some datasets can exhibit gaps as large as $10$ percentage points. These algorithms struggle to scale to datasets with hundreds or thousands of features or to deeper trees. It seems that we should return to greedy methods for larger-scale problems, but this would come at a loss of performance. Ideally, we should leverage greed only when it does not significantly deviate from optimality and use dynamic programming otherwise. Dynamic programming approaches build trees recursively, downward from the root. Problems farther from the root contain fewer samples and produce fewer splits. As we show, greedy splits near the root sacrifice performance, while greedy splits near the leaves produce performance close to the optimal. This suggests that we can tolerate less precision on problems close to leaves than on problems closer to the root – and that full optimization on those problems closer to the leaves yields only marginal returns relative to greedy, since we only have a few splits remaining. This has enormous implications, since the number of candidate trees increases exponentially with increases in depth; using greedy splitting closer to the leaves of the tree massively reduces the search space.
@@
Our contributions are as follows.
- We develop a family of decision tree algorithms that scale with the dataset size and number of features comparably to standard greedy algorithms but produce trees that are as accurate and sparse as optimal ones (e.g., Lin et al., 2020).
@@ 2 Related Work
Among the many methods for globally optimizing trees, Branch-and-bound approaches with dynamic programming are state of the art for scalability, because they exploit the structure of decision trees (Costa & Pedreira, 2023; Lin et al., 2020; Demirović et al., 2022; McTavish et al., 2022; Aglin et al., 2020). While many other methods exist for optimizing trees, such as MIP solvers (Bertsimas & Dunn, 2017; Verwer & Zhang, 2019), we focus our discussion and comparison of globally optimal decision tree methods on the currently fastest types of approaches – dynamic programming with branch and bound (DPBnB). These approaches search through the space of decision trees while tracking lower and upper bounds of the overall objective at each split to reduce the search space. They can find optimal trees on medium-sized datasets with tens of features and shallow maximum tree depths (Sullivan et al., 2024; Aglin et al., 2020; Lin et al., 2020; Demirović et al., 2022). Aglin et al. (2020) uses a DPBnB method with advanced caching techniques to find optimal decision trees, though it does not explicitly optimize for sparsity. In contrast, Lin et al. (2020); Hu et al. (2019) use a DPBnB approach to find a tree that optimizes a weighted combination of empirical risk and sparsity, defined by the number of leaves in the tree. McTavish et al. (2022) further enhances this approach by incorporating smart guessing strategies to construct tighter lower bounds for DPBnB, resulting in computational speedups. Demirović et al. (2022) extends the work of Aglin et al. (2020) by focusing on finding the optimal tree with a hard constraint on the number of permissible nodes, using advanced caching techniques and an optimized depth-2 decision tree solver. Mazumder et al. (2022) addresses continuous features by defining lower and upper bounds based on quantiles of feature distributions. However, their method is applicable only to shallow optimal trees with depth $\leq 3$, limiting its utility in scenarios with higher-order feature interactions.
@@
Several other approaches are compatible with branch and bound techniques. Blanc et al. (2024) seek to bridge the gap between greedy and optimal decision trees by selecting a fixed subset of the top $k$ feature splits for each sub-problem. However, this framework does not explicitly account for sparsity. Further, the method is limited by using a global setting for search precision: the approach considers the same number of candidate splits at each subproblem. As we show in our experiments, there is merit to tailoring the level of search precision to parts of the search space where it is most needed. The Blossom algorithm (Demirović et al., 2023) traverses a branch and bound dependency graph structure while using greedy heuristics to guide the search order. Relative to our approach, this algorithm optimizes from the bottom up, starting with greedy splits at each level, then optimizing the splits furthest from the root first. This choice guarantees eventual optimality while giving anytime behavior, but misses out on leveraging the property motivating this work – that greedy splits are most detrimental near the top of the tree. Like the approach of Blanc et al. (2024), Blossom also does not account for sparsity.
There are a few methods that use probabilistic search techniques to optimize trees. Sullivan et al. (2024) take a Bayesian approach, finding the maximum-a-posteriori tree by optimizing over an AND/OR graph, akin to the graph used in earlier branch-and-bound methods like that of Lin et al. (2020). Although their method demonstrates strong performance, their experimental results reveal that it is not responsive to sparsity-inducing hyperparameters – accordingly, we found in our experiments that the method struggles to optimize for sparsity.
@@ 3 Preliminaries
In Section 3, we use Rashomon sets to investigate properties of near-optimal trees.
@@
4 Algorithm Motivation
A key motivating property of SPLIT is that we can find high quality trees even when splitting greedily far from the root of the tree. To support this intuition, we empirically investigate how frequently near optimal trees behave greedily far from the root. To do so, we first generate the Rashomon set of decision trees for various values of sparsity penalty $\lambda$ and Rashomon bound $\epsilon$. Let $T\in\mathcal{R}(D,\lambda,\epsilon,d)$ be a tree in the Rashomon set, and let $n\in T$ be any node in $T$. Then, we compute the fraction of all nodes at a given level $\ell\leq d$ (where level 00 corresponds to the root) that were greedy (by which we mean that the split at this node in the tree is optimal with respect to information gain). This corresponds to the following proportion:
$\frac{\sum\limits_{T\in\mathcal{R}(D,\lambda,\epsilon,d)}\sum\limits_{\textrm{$n$}\in T}\mathbbm{1}[\textrm{$n$ is greedy $\land$ level$(n)=\ell$}]}{\sum\limits_{T\in\mathcal{R}(D,\lambda,\epsilon,d)}\sum\limits_{\textrm{$n$}\in T}\mathbbm{1}[\textrm{level}(n)=\ell]}.$ (3)
Figure 2 shows the results of this investigation for $6$ different datasets for different values of $\epsilon$ and $\lambda$. We note that there is a general increase in percentage of greedy splits as one goes deeper in the tree.
Figure 2: A heatmap of the proportion of splits of trees in the Rashomon set that are greedy, stratified by level, for different ($\lambda,\epsilon$) combinations. Only $4$ levels are shown as the $5^{th}$ level corresponds to the leaf. The greyed out regions in the bottom right of a plot represent ($\lambda,\epsilon$) for which the Rashomon set did not contain any trees of that depth. Generally, as we approach the leaves, the proportion of splits appearing in $\epsilon$-optimal trees become increasingly greedy. This is especially noticeable for the Netherlands, Covertype, and COMPAS datasets.
Additional motivating empirical results for using greedy splits far from the root of the tree are provided in Appendix A.2.
@@ 5.1 SParse Lookahead for Interpretable Trees (SPLIT)
$\begin{cases}\lambda+\min\Bigg{\{}\frac{|D^{-}|}{|D|},\frac{|D^{+}|}{|D|}\Bigg{\}}\ \ \hskip 44.10185pt\text{if $d^{\prime}=0$}\\ \\ \begin{aligned} &\min\Bigg{\{}\lambda+\frac{|D^{-}|}{|D|},\lambda+\frac{|D^{+}|}{|D|},\\ &\min_{f\in\mathcal{F}}\Big{\{}L\Big{(}T_{g}\big{(}D(f),d^{\prime},\lambda\big{)}\Big{)}+L\Big{(}T_{g}\big{(}D(\bar{f}),d^{\prime},\lambda\big{)}\Big{)}\Big{\}}\Bigg{\}}\end{aligned}\ \ \\ \hskip 142.26378pt\text{if $d^{\prime}=d-d_{l}$}\\ \begin{aligned} &\min\Bigg{\{}\lambda+\frac{|D^{-}|}{|D|},\lambda+\frac{|D^{+}|}{|D|},\\ &\min_{f\in\mathcal{F}}\Big{\{}\mathcal{L}\Big{(}D(f),d^{\prime}-1,\lambda\Big{)}+\mathcal{L}\Big{(}D(\bar{f}),d^{\prime}-1,\lambda\Big{)}\Big{\}}\Bigg{\}}\end{aligned}\ \ \\ \hskip 142.26378pt\text{if $d^{\prime}>d-d_{l}$.}\end{cases}$ (4)
@@ 6 Theoretical Analysis of Runtime and Optimality
Corollaries 6.2 and 6.3 show that, compared to globally optimal approaches, we see substantial improvements in runtime when lookahead depth is around half the global search depth.
@@ 7 Experiments
Figure 4: A comparison between the performance of our algorithms and competitors (depth budget $5$, lookahead depth $2$). The red box in the upper plot illustrates the region containing sparse and accurate models. The lower plots show the test loss vs training time for models in the red box. SPLIT and LicketySPLIT consistently lie on the bottom left of the test loss-sparsity frontier, with runtimes orders of magnitude faster than many competitors. Our algorithms also offer the ideal compromise between runtime and loss. All metrics are averaged over $3$ test-train splits.
@@ 8 Conclusion
Acknowledgements
We acknowledge funding from the National Institutes of Health under 5R01-DA054994, the National Science Foundation under award NSF 2147061, and through the Department of Energy under grant DE-SC0023194. We thank Srikar Katta for helpful discussions and feedback throughout this project.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1.2 What about depth $4$ trees?
In this section, we perform the same evaluation as above, but with depth $4$ trees. We set the lookahead depth as $2$.
Figure 6: A performance comparison between our algorithm and those in literature. Depth $4$ – Lookahead depth $2$.
@@ A.1.3 What about depth $6$ trees?
Figure 7: A performance comparison between our algorithm and those in literature. Depth $6$ – Lookahead depth $2$.
A.2 Many Near-Optimal Trees Exhibit Monotonically Decreasing Optimality Gaps Closer to Leaves
@@
$\beta(T,D,\ell)=\frac{\sum\limits_{\textrm{$t$}\in T}\delta(D_{t},t)\mathbbm{1}[\textrm{$t$ is rooted at level $\ell$}]}{{\sum\limits_{\textrm{$t$}\in T}\mathbbm{1}[\textrm{$t$ is rooted at level $\ell$}]}}.$ (6)
@@
A.3 Miscellaneous Properties of SPLIT
A.3.1 Which Lookahead Depth Should I Use?
@@
A.3.2 Are SPLIT Trees in the Rashomon Set?
@@
A.3.3 SPLIT with Optimality Preserving Discretization
@@
A.3.4 What is the performance gap between GOSDT post-processing for SPLIT / LicketySPLIT and purely greedy post-processing?
@@
A.4 SPLIT and LicketySPLIT Scaling Experiments
@@
A.5 Rashomon Importance Distribution Under RESPLIT vs TreeFARMS: Threshold Guessing
@@
A.6 Rashomon Importance Distribution Under RESPLIT: Quantile Binarization
@@
A.7 Experimental Setup
A.7.1 Datasets
@@ A.7.2 Preprocessing
    • For version $1$, we chose $n_{est}$ for each dataset such that the resulting binarized dataset has between $40$-$100$ features. This is the version used for experiments in Figures 4, 5, 6, 7 when we compare SPLIT / LicketySPLIT with other datasets.
@@
A.7.3 Details of Comparative Experiments for SPLIT and LicketySPLIT
@@
A Note on Comparative Experiments with Blossom (Demirović et al., 2023)
We briefly compare SPLIT with Blossom, an anytime decision tree algorithm incorporates greedy heuristics to guide search order, albeit in a bottom up manner. To our understanding, Blossom has no hyperparameters we can tune (except depth budget, min size, and min depth), which limits its flexibility in adapting to various datasets.
Dataset Runtime (s) Test Loss # Leaves
Blossom LicketySPLIT Blossom LicketySPLIT Blossom LicketySPLIT
compas 0.442 [0.381, 0.476] 0.334 [0.332, 0.336] 0.314 [0.303, 0.323] 0.317 [0.305, 0.329] 32.0 [32.0, 32.0] 8.0 [8.0, 8.0]
adult 16.223 [15.556, 16.744] 1.459 [1.453, 1.465] 0.177 [0.175, 0.179] 0.177 [0.173, 0.181] 32.0 [32.0, 32.0] 7.3 [6.4, 8.3]
netherlands 6.161 [6.128, 6.194] 0.627 [0.622, 0.632] 0.287 [0.282, 0.291] 0.292 [0.288, 0.296] 32.0 [32.0, 32.0] 7.7 [6.7, 8.6]
heloc 27.179 [26.810, 27.548] 0.510 [0.502, 0.518] 0.286 [0.281, 0.291] 0.294 [0.286, 0.303] 32.0 [32.0, 32.0] 7.0 [6.2, 7.8]
spambase 18.334 [18.200, 18.478] 0.487 [0.482, 0.492] 0.090 [0.085, 0.094] 0.087 [0.085, 0.088] 32.0 [32.0, 32.0] 13.7 [13.2, 14.1]
bike 158.744 [157.138, 159.964] 1.679 [1.633, 1.725] 0.112 [0.108, 0.115] 0.121 [0.117, 0.125] 32.0 [32.0, 32.0] 12.3 [11.9, 12.8]
bank 11.452 [11.053, 11.673] 0.353 [0.346, 0.360] 0.105 [0.097, 0.112] 0.103 [0.098, 0.108] 32.0 [32.0, 32.0] 9.0 [8.2, 9.8]
hypothyroid 2.799 [2.699, 2.909] 0.206 [0.198, 0.214] 0.004 [0.002, 0.006] 0.001 [0.000, 0.002] 17.8 [15.4, 20.2] 6.0 [6.0, 6.0]
covertype 13.617 [13.244, 13.861] 11.864 [11.577, 12.151] 0.237 [0.236, 0.238] 0.242 [0.241, 0.243] 32.0 [32.0, 32.0] 5.0 [5.0, 5.0]
Table 5: Comparison of Blossom and LicketySPLIT across datasets when Blossom is allowed to finish. We show the LicketySPLIT configuration (after grid-search across $\lambda$) that yielded the best test loss on that dataset (averaged across $5$ trials with depth budget $5$). Only the mean is bolded when LicketySPLIT performs better. Values are reported as mean [lower, upper] (indicating 95% confidence intervals).
From Table 5, we see that, despite being allowed to run to completion (taking over 10×10$\times$10 × longer in many cases), Blossom often underperforms LicketySPLIT in test loss. Furthermore, it is much less sparse than LicketySPLIT, having over 4×4$\times$4 × as many leaves for similar test performances.
We ran another experiment to examine Blossom’s anytime performance. In order to facilitate a fair comparison, we made Blossom run for approximately the same amount of time as LicketySPLIT (i.e. generally $\sim 1$ second) on a given dataset. Table 6 shows the best performing tree found by LicketySPLIT (found by varying $\lambda$ and computing the resulting test loss) for each dataset compared with a tree found by Blossom (depth $5$). We see that, given comparable runtimes, LicketySPLIT often achieves lower test loss with much fewer leaves compared to Blossom (which mostly branches out to $32$ leaves).
Dataset Runtime (s) Test Loss # Leaves
Blossom LicketySPLIT Blossom LicketySPLIT Blossom LicketySPLIT
compas 0.436 [0.404, 0.454] 0.334 [0.332, 0.336] 0.314 [0.303, 0.323] 0.317 [0.305, 0.329] 32.000 [32.000, 32.000] 8.000 [8.000, 8.000]
heloc 0.996 [0.992, 1.000] 0.510 [0.502, 0.518] 0.285 [0.281, 0.289] 0.294 [0.286, 0.303] 32.000 [32.000, 32.000] 7.000 [6.184, 7.816]
bike 1.896 [1.882, 1.908] 1.679 [1.633, 1.725] 0.128 [0.125, 0.133] 0.121 [0.117, 0.125] 30.000 [30.000, 30.000] 12.333 [11.862, 12.805]
covertype 14.297 [13.846, 14.604] 11.864 [11.577, 12.151] 0.237 [0.236, 0.238] 0.242 [0.241, 0.243] 32.000 [32.000, 32.000] 5.000 [5.000, 5.000]
adult 1.722 [1.717, 1.727] 1.459 [1.453, 1.465] 0.176 [0.174, 0.179] 0.176 [0.173, 0.181] 32.000 [32.000, 32.000] 7.333 [6.390, 8.276]
netherlands 1.238 [1.230, 1.246] 0.627 [0.622, 0.632] 0.284 [0.280, 0.287] 0.292 [0.288, 0.296] 32.000 [32.000, 32.000] 7.667 [6.724, 8.610]
bank 0.771 [0.765, 0.777] 0.353 [0.346, 0.360] 0.107 [0.099, 0.112] 0.103 [0.098, 0.108] 32.000 [32.000, 32.000] 9.000 [8.184, 9.816]
hypothyroid 0.649 [0.635, 0.663] 0.206 [0.198, 0.214] 0.004 [0.002, 0.006] 0.001 [0.000, 0.002] 18.800 [16.400, 21.200] 6.000 [6.000, 6.000]
spambase 0.832 [0.811, 0.856] 0.487 [0.482, 0.492] 0.091 [0.085, 0.097] 0.087 [0.085, 0.088] 32.000 [32.000, 32.000] 13.667 [13.196, 14.138]
Table 6: Comparison of Blossom and LicketySPLIT in an anytime setting (i.e. Blossom execution is stopped around the same time as LicketySPLIT) . We show the LicketySPLIT configuration (after grid-search across $\lambda$) that yielded the best test loss on that dataset (averaged across $5$ trials with depth budget $5$). Only the mean is bolded where LicketySPLIT outperforms Blossom. Values are reported as mean [lower, upper] (indicating $95\%$ confidence intervals). Note that the Blossom algorithm is not able to explicitly account for sparsity, hence it always returns fully grown trees up to a given depth.
A.7.4 Description of Machines Used